# Understanding via Gaze: Gaze-based Task Decomposition for Imitation Learning of Robot Manipulation

### v1

@@ Abstract
In imitation learning for robotic manipulation, decomposing object manipulation tasks into multiple semantic actions is essential. This decomposition enables the reuse of learned skills in varying contexts and the combination of acquired skills to perform novel tasks, rather than merely replicating demonstrated motions. Gaze, an evolutionary tool for understanding ongoing events, plays a critical role in human object manipulation, where it strongly correlates with motion planning. In this study, we propose a simple yet robust task decomposition method based on gaze transitions. We hypothesize that an imitation agent’s gaze control—fixating on specific landmarks and transitioning between them—naturally segments demonstrated manipulations into sub-tasks. Notably, our method achieves consistent task decomposition across all demonstrations, which is desirable in contexts such as machine learning. Using teleoperation—a common modality in imitation learning for robotic manipulation—we collected demonstration data for various tasks, applied our segmentation method, and evaluated the characteristics and consistency of the resulting sub-tasks. Furthermore, through extensive testing across a wide range of hyperparameter variations, we demonstrated that the proposed method possesses the robustness necessary for application to different robotic systems.
@@ I INTRODUCTION
Imitation stands as a uniquely promising approach for enabling robots to function as autonomous agents capable of acquiring object manipulation skills that align with human intentions. In fact, not only robots but also humans learn how they can and should handle various objects by observing expert demonstrations—such as young children learning from adults, or novices learning from skilled practitioners [1]. However, simply replicating demonstrated motions does not allow an agent to reuse learned skills in slightly altered contexts, nor to combine previously acquired motions into novel object manipulations.
To move beyond such a rote imitation, it is crucial that imitative agents gain a deeper understanding of demonstrated object manipulations, which requires autonomously segmenting a sequence of demonstrated motions into semantically meaningful units. The importance of such uniform task decomposition as the basis of action understanding is suggested by human studies—such as those by Newtson et al.—showing remarkable consistency among observers when segmenting video-recorded human actions [2].
@@
Figure 1: Gaze itself possesses an inherent structure in object manipulation, which enables task decomposition by detecting transitions between gaze landmarks during imitation. (a) Gaze transition of a human teleoperating a robot. (b) Gaze transition of a human making a sandwich [3]. (c) Overview of our proposed method.
@@
How, then, can we achieve autonomous and practical task decomposition of object manipulations? As Yarbus famously stated, “Eye movement reflects the human thought processes; so the observer’s thought may be followed to some extent from records of eye movement,” suggesting that eye gaze serves as a naturally evolved tool for understanding ongoing events [4]. In particular, it is well-established that during human object manipulations, individuals repeatedly fixate their gaze on task-relevant landmarks for certain periods and then move on to the next (Fig. 1), and that these gaze transition behaviors are tightly coupled with motor planning processes [5, 6, 7]. Furthermore, the work of Fathi et al. demonstrates that visual features located in the vicinity of a subject’s gaze during daily activities correlate with the verbal descriptions of these actions [3].
Taken together, these insights suggest that focusing on gaze position and the associated visual features offers a pathway to understanding—and thus segmenting—object manipulations. In this paper, we show that a simple approach of detecting transitions between gaze landmarks during imitation enables autonomous, general, and robust task decomposition of object manipulation sequences (Fig. 1). Here, although it would be preferable for the robot to autonomously control its own gaze during imitation, acquiring a general-purpose gaze-control policy remains challenging. Therefore, following the method proposed in [8], we measure the expert’s gaze during imitation (Fig. 1) and use it as a stand-in for the robot’s gaze. In other words, we assume an ideal scenario of joint attention, where the imitating agent and the human demonstrator share identical gaze transitions throughout the imitation process. This substitution is made feasible by the fact that the expert demonstrations are generated through robot teleoperation by a human. Under teleoperation, object manipulation typically relies solely on visual feedback, placing the human teleoperator under the same conditions as the imitating agent, which must interpret visual information to understand the demonstrated object manipulations.
In summary, the main contributions of this paper are follows:
  • We propose a practical task decomposition method that decomposes object manipulations in demonstrations into sub-tasks in a simple yet robust manner, by detecting the teleoperator’s gaze transitions during demonstrations.
  • We show that an imitation agent’s gaze control can simultaneously achieves task decomposition, underscoring the importance of developing autonomous gaze control capabilities in robotic object manipulation.
@@ II-A Robot Manipulation Imitation Using Task Decomposition
Many approaches have been proposed that utilize temporal task decomposition for imitation learning in object manipulation tasks. A common strategy for enabling robustness to long-horizon tasks is to employ goal-conditioned policy models [9, 10]. Furthermore, recent years have seen an increase in language-conditioned models aimed at multi-task capabilities [11, 12, 13, 14]. These goal- or language-conditioned methods implicitly assume that object manipulations in the given demonstrations are segmented into individual sub-tasks, and it has also been reported that providing language instructions per sub-task during inference improves task success rates [15].
However, in most existing work, the segmentation of these demonstrations is performed manually. Automating this process, as proposed in our research, can enhance robot autonomy and reduce annotation costs.
@@ II-B Conventional Task Decomposition Methods
A variety of methods have been proposed for temporally segmenting actions—including object manipulation—over time. Nonetheless, few are suited specifically for segmenting demonstrations of robot object manipulation.
Among the existing methods are those that segment actions based on hand or body trajectories and velocities [16, 17, 18]. Such methods require body tracking and are not robust to trajectory variations. In contrast, our proposed approach does not rely on directly measuring or leveraging the motion itself. Other methods detect segmentation points through changes in the relationship between objects and the hand [19, 20], but they demand advanced image processing techniques, such as object or hand segmentation, detection, and tracking. Although further investigation is required to determine which type of segmentation is more beneficial for imitation learning, there is a qualitative difference between the segmentation results of these methods and those produced by our approach. Notably, our method does not necessarily cut actions at the moment of contact.
With the advent of deep learning, there has been a rise in end-to-end approaches that take images as input and directly output action labels [21, 22, 23, 24]. However, applying these methods to previously unseen tasks requires additional data and annotations, and they do not guarantee uniform segmentation. By contrast, our method can perform task decomposition without any supervision, as long as gaze data is available. Although there have been a few studies focusing on gaze—particularly on image features around the gaze position [25, 3]—our approach explicitly ties the timing of gaze transitions to the timing of sub-task changes, achieving a simple and uniform decomposition.
@@ III METHOD
In this section, we first formulate task decomposition of object manipulation based on gaze transitions. We then detail our implementation procedure. Specifically, we begin by explaining how gaze transitions are detected from time-series gaze data, and subsequently describe the refinement process used to achieve consistent segmentation across the demonstration dataset.
@@ III-A Gaze-based Task Decomposition
We consider a scenario in which an agent that has acquired gaze control imitates object manipulation using only visual information. Let $\{g_{t}\}_{t=0}^{T}$ denote the time-series of the agent’s gaze positions, and $\{f_{t}\}_{t=0}^{T}$ the time-series of the visual features observed around those gaze positions.
When the object manipulation is decomposed into a set of sub-tasks $\{\tau_{k}\}_{k=0}^{K}$, the objective in this section is to find the $K-1$ time steps at which the sub-tasks switch (i.e. $\{t\mid k(t+1)=k(t)+1\}$).
In human object manipulation, it is well-established that each task has a specific set of gaze landmarks $\{c_{i}\}_{i=0}^{N}$, and that humans fix their gaze on each landmark for a certain duration, then sequentially transition to the next landmark [6, 7, 5]. Here, the landmark index $i$ is arranged according to the order in which landmarks are fixated from the start of the task (duplicates may occur).
In this study, we hypothesize that the time steps at which the index of the fixated gaze landmark changes, $\{t\mid i(t+1)=i(t)+1\}$, corresponds to the time steps at which the sub-tasks switch, $\{t\mid k(t+1)=k(t)+1\}$. In other words, we realize task decomposition by detecting transitions of gaze landmarks.
Let $p^{i(t)}_{t}$ be the location of the landmark $c_{i(t)}$ in the agent’s field of view at time step $t$. Then, the gaze position $g_{t}$ and the visual feature $f_{t}$ at that gaze position can be written as follows:
$g_{t}=p\bigl{(}c_{i=i(t)},t\bigr{)}+\epsilon_{t},$ (1)
$f_{t}=f\bigl{(}g_{t}\bigr{)}+\zeta_{t}=f\bigl{(}p(c_{i=i(t)},t)+\epsilon_{t}\bigr{)}+\zeta_{t}.$ (2)
where $\epsilon_{t}$ and $\zeta_{t}$ represent noise in the gaze position and perturbations in the visual features caused, for example, by hand movements, respectively, and $f(\cdot)$ is the visual feature extractor at the given gaze position.
In other words, by leveraging the relationships in Eqs. (1) and (1), we can detect the transitions of gaze landmarks from the gaze information $\{g_{t}\}$ and $\{f_{t}\}$, and thereby achieve task decomposition of the demonstrated object manipulation.
@@ III-B Gaze Transition Detection
We now describe the implementation details for segmenting demonstration tasks based on gaze transitions. As discussed in the previous section, an algorithm for detecting the time steps at which gaze landmarks switch, using time-series gaze data, is critical for task decomposition.
@@ III-B1 Median Filtering
Gaze data typically contain noise. Even when the agent is fixating on a stationary gaze landmark, the recorded gaze positions often exhibit fluctuations. Moreover, the agent may occasionally shift its gaze for a short period to a different landmark during object manipulation. To handle such noise in a simple yet effective manner, we adopted a median-filtering approach. Specifically, for the agent’s gaze position $g_{t}$ we applied the process shown in Eq. (3) with a window size $w$ as a hyperparameter, and used the filtered gaze data $\hat{g}_{t}$ in all subsequent processing:
@@ III-B2 Visual Features Around the Gaze
Gaze transitions cause not only positional changes in the gaze but also variations in the visual features around the gaze. To detect these transitions from visual features, it is necessary to employ a feature extractor based on high-quality representation learning. For each time step, we cropped an $b\times b$ region centered at the filtered gaze position $\hat{g}_{t}$ from both the left and right images and fed these cropped images into the pre-trained CLIP model [26] to extract their feature vectors $f_{t}\in\mathbb{R}^{512\times 2}$. This time-series of feature vectors, along with the gaze position, was used for gaze transition detection.
@@ III-B3 Change Detection
For the obtained gaze position $\hat{g}_{t}$ and visual feature $f_{t}$, we computed the difference from the previous time step to detect abrupt changes in the time-series data caused by gaze transitions:
@@
Here, $s^{pos}_{t}$ and $s^{feat}_{t}$ represent the change scores for the gaze position and the visual features, respectively. Specifically, $s^{pos}_{t}$ is the Euclidean distance between the two 4-dimensional vectors representing the left and right gaze positions. Meanwhile, $s^{feat}_{t}$ is derived from the inner products between the corresponding gaze feature vectors in consecutive time steps, normalized to ensure non-negativity, and then averaged for the left and right features.
@@
We introduced the thresholds $\theta_{pos}$ and $\theta_{feat}$ as hyperparameters for these two change scores, and identified any time step at which both change scores exceeded their respective thresholds as a change point $d_{i}\in D$.
III-C Refinement Based on Segmentation Uniformity
From a machine learning perspective—particularly one that relies on multiple demonstrations—it is desirable for all demonstrations of the same task to undergo a uniform task decomposition. For example, in a pick-and-place task, mixing one demonstration segmented into “pick” and “place” with another in which these two motions are grouped into a single segment is not ideal for building a consistent training dataset. Therefore, in this study, we introduce a refinement procedure, $D\rightarrow D^{*}$, for each demonstration to ensure that demonstrations of the same task share as consistent a number of sub-tasks as possible.
@@ III-C Refinement Based on Segmentation Uniformity
First, we enumerate the number of change points $\lvert D\rvert$ obtained via the gaze transition detection (Section III-B) for all demonstrations and denote the most frequent value by $s$. We assume $s$ to be the number of sub-tasks for the task and adjust the thresholds $\theta_{pos}$ and $\theta_{feat}$ so that the number of segments in each demonstration matches $s$. Specifically, for any demonstration whose segmented count is smaller than $s$, we reduce both thresholds by 1% at each iteration; conversely, for a count larger than $s$, we increase them by 1%. If no threshold values are found that yield $s$ segments within a maximum number of iterations, that demonstration is treated as a detection failure and excluded from the training data.
@@
Finally, the refined set of change points $D^{*}$ obtained by this process is taken as the sub-task boundary for each demonstration, enabling us to divide as many demonstrations as possible into a consistent number of sub-tasks.

### v2

# Gaze-Based Task Decomposition for Robot Manipulation in Imitation Learning

@@ Abstract
In imitation learning for robotic manipulation, decomposing object manipulation tasks into multiple sub-tasks is essential. This decomposition enables the reuse of learned skills in varying contexts and the combination of acquired skills to perform novel tasks, rather than merely replicating demonstrated motions. Gaze plays a critical role in human object manipulation, where it is strongly correlated with hand movements. We hypothesize that an imitating agent’s gaze control—fixating on specific landmarks and transitioning between them—simultaneously segments demonstrated manipulations into sub-tasks. In this study, we propose a simple yet robust task decomposition method based on gaze transitions. The method leverages teleoperation, a common modality in robotic manipulation for collecting demonstrations, in which a human operator’s gaze is measured and used for task decomposition as a substitute for an imitating agent’s gaze. Notably, our method achieves consistent task decomposition across all demonstrations for each task, which is desirable in contexts such as machine learning. We applied this method to demonstrations of various tasks and evaluated the characteristics and consistency of the resulting sub-tasks. Furthermore, through extensive testing across a wide range of hyperparameter variations, we demonstrated that the proposed method possesses the robustness necessary for application to different robotic systems.
@@ I INTRODUCTION
Figure 1: Gaze itself possesses an inherent structure in object manipulation, which enables task decomposition by detecting transitions between gaze landmarks during demonstration. (a) Gaze transition of a human teleoperating a robot. (b) Overview of our proposed method.
Imitation stands as a uniquely promising approach for enabling robots to function as autonomous agents capable of acquiring object manipulation skills that align with human intentions. Here, the ability for an imitating agent to automatically decompose a demonstrated manipulation into multiple sub-tasks is instrumental for executing long-horizon tasks and for reusing or combining learned skills.
In this context, a method is required that can segment demonstrations of various tasks into sub-tasks without the need for additional training or diverse data collection. In addition, recent imitation learning methods based on deep learning typically require hundreds of demonstrations for each task, and it is desirable to consistently decompose all of these demonstrations into the same set of sub-tasks. However, few methods have been proposed for such task decomposition, and none of them are able to perform it in a consistent manner. As a result, demonstrations are often manually segmented into sub-tasks [1], otherwise the goal state, which conditions a policy model, is usually substituted with a state a certain number of steps ahead [2, 3].
@@
In this study, we focus on gaze information to achieve consistent task decomposition without the need for additional training or diverse data collection. As Yarbus stated, “Eye movement reflects the human thought processes; so the observer’s thought may be followed to some extent from records of eye movement” [4], suggesting that gaze is a naturally evolved tool for understanding ongoing events. In particular, it is well established that during human object manipulations, individuals repeatedly fixate their gaze on task-relevant landmarks for certain periods and then move on to the next (Fig. 1), and that these gaze transition behaviors are tightly coupled with the motor planning processes of hands [5, 6, 7]. These insights suggest that leveraging an imitating agent’s gaze information, including its position and associated visual features, provides a pathway for task decomposition in demonstrated object manipulations.
In this paper, we show that a simple approach of detecting transitions between gaze landmarks during imitation enables robust and consistent task decomposition of demonstrations (Fig. 1). In the current imitation learning framework, it is common to collect demonstrations via human teleoperation. Several approaches have been proposed that simultaneously measure the teleoperator’s gaze alongside motion recording and use it as a substitute for an imitating agent’s gaze. [8, 9]. These methods improved both the robustness and dexterity of the imitated actions with minimal modifications to the conventional teleoperation system. Similar to these methods, we record the teleoperator’s gaze information (Fig. 1) and use it for task decomposition. Our proposed method segments demonstrations at the operator’s gaze transition timings, identified using a threshold-based approach that detects abrupt changes in gaze position and visual features around the gaze point. To ensure consistent task decomposition across all demonstrations of the same task, we also develop a refinement process that automatically adjusts these thresholds for each demonstration.
We apply our method to demonstration datasets of three tasks collected for imitation learning and evaluated the characteristics and consistency of the resulting sub-tasks. Furthermore, we investigate how the proposed refinement process improves the consistency of task decomposition and the robustness to variations in hyperparameters.
@@ II-A Robot Manipulation Imitation Using Task Decomposition
Many approaches have been proposed that utilize temporal task decomposition for imitation learning in robot manipulation. A common strategy for enhancing robustness in long-horizon tasks is to employ goal-conditioned policy models [1, 2]. In these methods, the task is decomposed into multiple sub-tasks, and the goal states of each sub-task are conditioned to the model stepwise. Furthermore, recent years have seen an increase in language-conditioned models aimed at multi-task capabilities [10, 11, 12, 13]. In datasets for language-conditioned models, sub-task-level instructions, such as ”pick an apple” and ”place on a plate”, are generally preferred over task-level instructions, such as ”move an apple to a plate” [14]. These goal- or language-conditioned methods implicitly assume that object manipulations in the given demonstrations are segmented into individual sub-tasks. However, in most existing work, the segmentation of these demonstrations is performed manually. In this context, our proposed method can reduce annotation costs by automating part of the data collection process, specifically by segmenting demonstrations.
@@ II-B Conventional Task Decomposition Methods
In the field of Action Segmentation, numerous methods have been proposed for the temporal segmentation of actions and motions involving object manipulation [15, 16, 17, 18, 19, 20, 21], including ones using human gaze information [22, 23]. However, most of these methods are primarily designed for human action recognition, and very few are suitable for task decomposition of robotic object manipulation. Moreover, because many of these approaches rely heavily on visual, data-driven techniques, applying them to a wide variety of robotic hardware requires the construction of new datasets tailored to each specific platform.
Although research on task decomposition in robotic object manipulation is limited, several studies have been proposed. Liang et al. optimize the transitions between the manipulating (tool) and the manipulated (target) objects, using these transition points to segment tasks; however, their method requires the 6D poses of all relevant objects [24]. In contrast, our proposed method does not require explicit modeling of the environment. Lee et al. reduce the dimensionality of whole-body trajectories using PCA and further model them with GMM, assigning each Gaussian component to a sub-task [25]. In contrast, our method does not utilize hand motion data, making it robust to variations in such movements. Zhu et al. employ neural networks (NNs) for latent-space clustering to achieve task decomposition; however, their approach requires training the NN for new tasks, and the number of sub-tasks is determined solely by the time step length of each segment [26]. In contrast, our method does not require training of NNs, thus enabling high-speed processing, and it determines the number of sub-tasks based on the number of gaze transitions, thus ensuring higher generality. Zhang et al. utilize latent vectors derived from pre-trained models specialized for object manipulation, such as R3M [27], and perform task decomposition based on the distances between these latent vectors [28]. Similar to our method, their approach employs detection-based task decomposition without requiring additional data or training. However, compared to distance metrics in latent vector space, gaze transitions have been extensively studied for their coordination with hand movement planning in object manipulation, indicating that gaze-based task decomposition offers greater reliability.
@@ III METHOD
In this section, we first formulate the task decomposition of object manipulation based on gaze transitions. We then detail our implementation procedure. Specifically, we begin by explaining how gaze transitions are detected from time-series gaze data, and subsequently describe the refinement process used to achieve consistent decomposition across the demonstration dataset.
@@ III-A Gaze-based Task Decomposition
We consider a scenario in which a human remotely controls a robot using only visual information to collect demonstrations of object manipulation for imitation learning. Multiple demonstrations (typically ranging from several tens to hundreds) are collected for each task, and all demonstrations of the same task share an identical set of sub-tasks. Let $\{g_{t}\}_{t=0}^{T}$ denote the time-series of the human teleoperator’s gaze positions, and $\{f_{t}\}_{t=0}^{T}$ the time-series of the visual features observed around those gaze positions. When the demonstrated task is decomposed into a set of sub-tasks $\{\tau_{k}\}_{k=0}^{K}$, the objective in this section is to determine the $K-1$ time steps at which the sub-tasks switch (i.e. $\{t\mid k(t+1)=k(t)+1\}$), using the gaze information $\{g_{t}\}$ and $\{f_{t}\}$.
In human object manipulation, it is well-established that each task has a specific set of gaze landmarks $\{l_{i}\}_{i=0}^{N}$, and that humans fix their gaze on each landmark for a certain duration, then sequentially transition to the next landmark [6, 7, 5]. Here, the landmark index $i$ is arranged according to the order in which landmarks are fixated from the start of the task (duplicates may occur).
In this study, we hypothesize that the time steps at which the index of the fixated gaze landmark changes, $\{t\mid i(t+1)=i(t)+1\}$, correspond to the time steps at which the sub-tasks switch, $\{t\mid k(t+1)=k(t)+1\}$. In other words, we realize task decomposition by detecting transitions of gaze landmarks.
Let $p^{i(t)}_{t}$ be the location of the landmark $c_{i(t)}$ in the teleoperator’s field of view at time step $t$. Then, the gaze position $g_{t}$ and the visual feature $f_{t}$ at that gaze position can be written as follows:
$g_{t}=p^{i(t)}_{t}+\epsilon_{t},$ (1)
$f_{t}=f\bigl{(}g_{t}\bigr{)}+\zeta_{t}=f\bigl{(}p^{i(t)}_{t}+\epsilon_{t}\bigr{)}+\zeta_{t},$ (2)
where $\epsilon_{t}$ and $\zeta_{t}$ represent noise in the gaze position and perturbations in the visual features caused, for example, by the end-effector movements, respectively, and $f(\cdot)$ is the visual feature extractor at the given gaze position.
In other words, by leveraging the relationships in Eqs. 1 and 2, we can detect the transitions of gaze landmarks $\{l_{i}\}$ from the gaze information $\{g_{t}\}$ and $\{f_{t}\}$, and thereby achieve task decomposition of the demonstrated object manipulation.
@@ III-B Gaze Transition Detection
We now describe the implementation details for segmenting demonstrations based on gaze transitions. As discussed in the previous section, an algorithm for detecting the time steps at which gaze landmarks switch, using time-series gaze data, is critical for the task decomposition.
@@ III-B1 Median Filtering
Gaze data typically contain noise. Even when the teleoperator is fixating on a stationary gaze landmark, the recorded gaze positions often exhibit fluctuations. Moreover, the teleoperator may occasionally shift its gaze for a short period to a different landmark during object manipulation. To handle such noise in a simple yet effective manner, we adopt a median-filtering approach. Specifically, for the teleoperator’s gaze position $g_{t}$, we apply the process shown in Eq. 3 with a window size $w$ as a hyperparameter, and use the filtered gaze data $\hat{g}_{t}$ in all subsequent processing:
@@ III-B2 Visual Features Around the Gaze
Gaze transitions cause not only positional changes in the gaze but also variations in the visual features around the gaze. We also consider detecting gaze transitions using visual features. To robustly detect changes in visual features, it is essential to employ a feature extractor based on high-quality representation learning. For each time step, we crop an $b\times b$ region centered at the filtered gaze position $\hat{g}_{t}$ from both the left and right images and feed these cropped images into the pre-trained CLIP model [29] to extract their feature vectors $f_{t}\in\mathbb{R}^{512\times 2}$. This time-series of feature vectors, along with the gaze position, is used for gaze transition detection in this study.
@@ III-B3 Change Detection
For the obtained gaze position $\hat{g}_{t}$ and visual feature $f_{t}$, we compute the difference from the previous time step to detect abrupt changes in the time-series data caused by gaze transitions:
@@
where $s^{pos}_{t}$ and $s^{feat}_{t}$ represent the change scores for the gaze position and the visual features, respectively. Specifically, $s^{pos}_{t}$ is the Euclidean distance between the two 4-dimensional vectors representing the left and right gaze positions. Meanwhile, $s^{feat}_{t}$ is derived from the inner products between the corresponding gaze feature vectors in consecutive time steps, normalized to ensure non-negativity, and then averaged for the left and right features.
@@
We introduce the thresholds $\theta_{pos}$ and $\theta_{feat}$ as hyperparameters for these two change scores, and identify any time step at which both change scores exceed their respective thresholds as a change point $c_{i}\in C$.
III-C Refinement Based on Decomposition Consistency
In recent machine learning–based imitation learning methods, particularly those relying on multiple demonstrations, all demonstrations of the same task typically comprise a consistent set of sub-tasks. In such scenarios, it becomes possible to further improve the consistency of task decomposition across all these demonstrations. Here, we introduce a refinement procedure, $C\rightarrow C^{*}$, to ensure that all demonstrations of the same task share as consistent a number of sub-tasks as possible. The algorithm for this process is summarized in Algorithm 1.
@@ III-C Refinement Based on Decomposition Consistency
First, we enumerate the number of change points $\lvert C\rvert$ obtained via the gaze transition detection (Section III-B) for every demonstration and denote the most frequent value by $s$. We assume $s$ to be the number of sub-tasks for the task and adjust the thresholds $\theta_{pos}$ and $\theta_{feat}$ so that the number of segments in each demonstration matches $s$. Specifically, for any demonstration whose segmented count is smaller than $s$, we reduce both thresholds by 1% at each iteration; conversely, for a count larger than $s$, we increase them by 1%. If no threshold values are found that yield $s$ segments within a maximum number of iterations, that demonstration is treated as a detection failure and excluded from the training data.
@@
Finally, the refined set of change points $C^{*}$ obtained by this process is taken as the sub-task boundary for each demonstration, enabling us to divide as many demonstrations as possible into a consistent number of sub-tasks.
Algorithm 1 Task Decomposition Refinement Process
1:Given: A set of demonstrations for the same task:
$D=\{(s_{t}^{\mathrm{pos}},\,s_{t}^{\mathrm{feat}})^{(n)}\}_{n=1}^{N},\quad t\in[0,\,T^{(n)}],$
where $N$ is the number of demonstrations and $T^{(n)}$ is the total time steps in the $n$-th demonstration.
2:Notation: Let the function $\textit{detect}(\cdot)$ perform change point detection:
$C\;\leftarrow\;\textit{detect}\bigl{(}\{s_{t}^{\mathrm{pos}}\},\,\{s_{t}^{\mathrm{feat}}\},\,\theta_{\mathrm{pos}},\,\theta_{\mathrm{feat}}\bigr{)}.$
3:Initialize: $\theta_{\mathrm{pos}},\theta_{\mathrm{feat}}\;\leftarrow\;\text{default values}$
4:for $n=1$ to $N$ do
5:     $C^{(n)}\;\leftarrow\;\textit{detect}\bigl{(}\{s_{t}^{\mathrm{pos}}\}^{(n)},\,\{s_{t}^{\mathrm{feat}}\}^{(n)},\,\theta_{\mathrm{pos}},\,\theta_{\mathrm{feat}}\bigr{)}$
6:end for
7:$s\;\leftarrow\;\text{mode}(\{|C^{(n)}|\}_{n=1}^{N})$
8:for $n=1$ to $N$ do
9:     while $\bigl{|}C^{(n)}\bigr{|}<s$ do
[…]
11:         $C^{(n)}\leftarrow\textit{detect}(\{s_{t}^{\mathrm{pos}}\}^{(n)},\{s_{t}^{\mathrm{feat}}\}^{(n)},\theta_{\mathrm{pos}},\theta_{\mathrm{feat}})$
12:     end while
13:     while $\bigl{|}C^{(n)}\bigr{|}>s$ do
[…]
15:         $C^{(n)}\leftarrow\textit{detect}(\{s_{t}^{\mathrm{pos}}\}^{(n)},\{s_{t}^{\mathrm{feat}}\}^{(n)},\theta_{\mathrm{pos}},\theta_{\mathrm{feat}})$
16:     end while
17:     $C^{*(n)}\;\leftarrow\;C^{(n)}$
18:end for
19:Return: $\{C^{*(n)}\}_{n=1}^{N}$